OpenStudy (anonymous):

1. Solve by factoring a. 3x^2-18x=0 b. x(3x-1)=2 c.(4a+1)(3a+1)=13 d.x^2-16x+48=0

OpenStudy (anonymous):

a. \[3x ^{2}-18x=0\] \[3x ^{2}=18\] \[x ^{2}=9\] \[x=3;x=-3\]

OpenStudy (anonymous):

3m^3-m^2-18m=0

OpenStudy (anonymous):

thank you guys for your help. i really appreciate it

OpenStudy (anonymous):

\[x(3x-1)=2\] \[3x ^{2}-x-2=0\] \[(3x+2)(x-1)\]

OpenStudy (anonymous):

c.\[(4a+1)(3a+1)=13\] \[12a ^{2}+4a+3a+1-13=0\] \[12a ^{2}+7a-12=0\] \[(4a-3)(3a+4)\]

OpenStudy (anonymous):

\[x ^{2}-16x+48=0\] \[(x-12)(x-4)\]

OpenStudy (anonymous):

ah yah @Euler271 is right sorry

OpenStudy (anonymous):

the mistake in a is actually more than just that. you don't consider when x = 0 you hypothetically divided by zero

OpenStudy (anonymous):

proper way is 3x(x-6) = 0 x = 0 or x = 6

OpenStudy (anonymous):

@abrown16 please consider @Euler271 answers for letter a mine is wrong :)

OpenStudy (anonymous):

ok @yeyenunez thank you